OpenStudy (anonymous):

Find the exact value of sin(arctan(2)).

OpenStudy (turingtest):

what triangle corresponds to arctan(2) ?

OpenStudy (anonymous):

it doesnt say but im assuming a right triangle

OpenStudy (turingtest):

yes, but which right triangle? if we have\[\tan^{-1}(2)=\theta\implies\tan\theta=\frac{\sin\theta}{\cos\theta}=2\]

OpenStudy (anonymous):

so we are then left with sin(sin 2/cos 2)

OpenStudy (turingtest):

no, we want the arctan of 2. not tan of 2

OpenStudy (turingtest):

remember the definitions of sine and cos

OpenStudy (anonymous):

so cos is on the top and sin on the bottom?

OpenStudy (turingtest):

why do you think that?

OpenStudy (turingtest):

do you know what arctan means?

OpenStudy (anonymous):

arctan is the opposite of tan

OpenStudy (turingtest):

in a sense... arctan(x) is the angle t at which tan(t)=x

OpenStudy (turingtest):

I think you have confused it with the cotangent

OpenStudy (anonymous):

tan= opposite/adjacent so that 2/1

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

then we have to find the hypoteneus so we can get the sin

OpenStudy (turingtest):

exactly :)

OpenStudy (anonymous):

thats 2/sqrt5

OpenStudy (turingtest):

yep :) nice job

OpenStudy (anonymous):

i see it makes since now. thank you!!

OpenStudy (turingtest):

welcome, happy to help